sandra uses two mobile cell phones (a Digicel and a Lime phone), to make her monthly calls. Sandra budget only $7,000 on telephone calls per month and the cost per minute for digicel and limeare $2.99 and $2.89 respectively.

write down the algebric equation for Sandra's budget constraint.
1
Expert's answer
2016-11-11T05:42:09-0500
Equation looks like
2.99x+2.89y=7,000
Where x is the quantity of minutes for digicel,
y is the quantity of minutes for limeare
Or in a more familiar form

Y=(7,000-2.99x)/2.89;
Y=2,42215-1.03x

Answer.
2.99x+2.89y=7,000